A coffee merchant has coffee beans that sell for $9 per poind and $12 per point. The two types are to be mixed to create 100lb of a mixture that will sell for $11.25 per pound.How much of each type of bean should be used in the mixture?
Answer by Earlsdon(6103)

Let x = the required number of pounds of $9.00/lb coffee beans and 100-x = the required number od $12.00/lb coffee beans. You can write the equation from this.
x($9.00) + (100-x)($12.00) = 100($11.25) Simplify and solve for x.
9x + 1200-12x = 1125 Combine the x-terms and subtract 1200 from both sides.
-3x = -75 Divide both sides by -3
x = 25
100-x = 75
The mixture should contain 25 lbs of $9.00/lb coffee beans and 75 lbs of $12.00/lb coffee beans.
Check:
25 lbs($9.00/lb) + 75 lbs($12.00/lb) = $225.00 + $900.00 = $1,125.00
100 lbs($11.25/lb) = $1,125.00